# Accelerating Materials Design via LLM-Guided Evolutionary Search

### v1

@@ 4.3 Impact of Surrogate Models
Figure 6: Hit Rate and Stability performance of LLEMA with and without surrogate model-based property prediction.
Figure 6 quantifies the effect of surrogate model–based property prediction on LLEMA’s performance. All experiments were run for 250 iterations with task constraints relaxed to encourage exploration, similar to the Section 4.2. To isolate the contribution of ML-based surrogates, we removed surrogate ML models like CGCNN and ALIGNN, and restricted the workflow to using only the Materials Project database for property annotations. Furthermore, we experimented with fewer iterations while relaxing task constraints for the wide-bandgap semiconductors dataset. Even in this setting, both hit rate and stability collapse to near-zero (<5<5%), as the evolutionary process cannot assign meaningful rewards without surrogate models to candidates for missing property annotations in the Materials Project API. The search then drifts toward trivial or repeated compounds instead of progressing toward novel solutions. By contrast, reintroducing surrogate predictors yields more than a sixfold improvement, raising hit rate and stability into the 25–30% range. The surrogate estimates for out-of-distribution candidates supply the evaluation signals needed to sustain exploration beyond the sparse coverage of existing datasets. These results demonstrate that surrogate models are indispensable for furnishing informative fitness signals under sparse supervision, preventing collapse, and enabling effective discovery in incomplete material spaces.
@@ 5 Related Work
Materials discovery has progressed from trial-and-error (Nagamatsu et al., 2001) to ab initio modeling (Jain et al., 2011; Hautier et al., 2012; Pyzer-Knapp et al., 2015), with DFT and high-throughput screening as standard tools. Machine learning further accelerates discovery via rapid property prediction (Xie & Grossman, 2018; Chen et al., 2019; Choudhary & DeCost, 2021) and generative design (Gebauer et al., 2019; Xie et al., 2021; Jiao et al., 2023), including multi-objective optimization (Gopakumar et al., 2018; Jablonka et al., 2021). Yet, these methods remain limited by data scarcity and poor transferability across domains. In contrast, LLMs leverage broad scientific knowledge for reasoning and hypothesis generation in data-poor regimes. Our work integrates such knowledge-driven reasoning with predictive oracles and domain rules to advance multi-objective materials synthesizability-aware materials design.
@@
Recent advances in generative models have shown their ability to generalize across diverse tasks using pre-trained knowledge and simple prompting strategies (Brown et al., 2020; Wei et al., 2022). However, their outputs are often unreliable or inconsistent (Madaan et al., 2024; Zhu et al., 2023), motivating the use of evolutionary optimization frameworks where LLMs act as generators and external evaluators guide selection and refinement (Lange et al., 2024; Lehman et al., 2023; Liu et al., 2024; Zheng et al., 2023). Such frameworks have been successfully applied in areas including code and prompt generation (Guo et al., 2023), mathematical optimization (Yang et al., 2024), program synthesis (Romera-Paredes et al., 2024), robotics reward design (Ma et al., 2023), and chemical discovery (Wang et al., 2024). We extend this line of work to real-world materials science challenges by incorporating chemistry-informed LLM evolution that enforces structural validity and physical plausibility, enabling principled discovery under multi-objective constraints.
@@
Early work applied LLM-based frameworks to literature mining, named-entity recognition, and property extraction (Gupta et al., 2022; 2024; Niyongabo Rubungo et al., 2025). Beyond text extraction, more recent methods use LLMs for hypothesis generation (Miret & Krishnan, 2024), synthesis route planning, and as reasoning engines in multi-agent systems (Zhang et al., 2024; Kang & Kim, 2024; Kumbhar et al., 2025). Systems such as MatAgent (Bazgir et al., 2025) and LLMatDesign (Jia et al., 2024) further combine LLM reasoning with property predictors and optimization loops. However, existing approaches often lack constraint enforcement, focus on narrow single-objective tasks, and tend to collapse into unguided search, limiting the generation of synthesizable, property-aligned candidates. LLEMA introduces an LLM-guided framework that integrates reasoning-driven candidate generation with surrogate predictors, domain-guided rules, and memory-based refinement, achieving a more effective balance between novelty, feasibility, and multi-objective property alignment.
@@ Acknowledgements
This research was partially supported by the U.S. National Science Foundation (NSF) under Grant No. 2416728. Saaketh Desai is supported in part by the Center for Integrated Nanotechnologies, an Office of Science user facility operated for the U.S. Department of Energy. This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title, and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan.
@@ Appendix A Autonomous Materials Discovery
The discovery of novel materials with tailored properties is fundamental to technological progress, contributing to a global materials industry generating approximately \$ 8 trillion in revenue in 2023 (McKinsey, 2024). The impact of tailored materials is particularly significant in critical domains such as energy storage Ling (2022), photovoltaics Solak & Irmak (2023), and microelectronics Refai-Ahmed et al. (2024), where advanced materials can enhance efficiency, reduce energy footprint, and improve sustainability. The grand challenge in materials discovery lies in navigating an enormous design space encompassing a vast number of material compositions and manufacturing techniques while achieving the target material structures and properties across multiple length and time scales (Oganov et al., 2019). Human intuition-based methods, along with traditional design of experiments and computational investigations, are slow and ineffective at screening this vast space of possible materials and synthesis conditions. Consequently, the field has witnessed a rapid emergence of data-driven discovery paradigms Agrawal & Choudhary (2016) and autonomous self-driving laboratories Abolhasani & Kumacheva (2023), integrating artificial intelligence, robotics, and high-throughput computation. Data-centric platforms such as the Materials Project (Jain et al., 2013) exemplify this shift, offering open high-throughput DFT databases that enable large-scale screening of candidate materials (Horton et al., 2025). Agent-based frameworks further advance this paradigm by autonomously navigating complex chemical spaces and identifying novel stable compounds (Montoya et al., 2020). Together, these systems mark a paradigm shift, transforming materials discovery from a process of human-guided trial and error to one of algorithmic intuition and autonomous exploration.
Appendix B Datasets and Benchmark
To evaluate multi-objective material discovery, we curate a diverse benchmark spanning 14 representative design tasks across various domains. Each task defines a distinct combination of physicochemical constraints that reflect practical design objectives in real-world materials engineering. Together, the benchmark captures the breadth of challenges faced in materials design, from optimizing performance–stability trade-offs to balancing mechanical, dielectric, and sustainability objectives. The resulting dataset assesses how well models can reason over complex, interdependent physical properties and generate synthesizable candidates under realistic constraints.
Wide-Bandgap Semiconductors.
SAW/BAW Acoustic Substrates.
High-$k$ Dielectrics.
Solid-State Electrolytes.
Solid electrolytes promise safe, high-energy batteries by replacing flammable liquid electrolytes. Candidates must be thermodynamically stable (formation energy $\leq-1.0$ eV/atom, energy above hull $\leq 0.1$ eV/atom), electronically insulating (band gap $\geq 2.0$ eV), and contain mobile species such as Li, Na, K, Mg, Ca, or Al. This task reflects the fundamental trade-off between chemical stability and ionic conductivity, which is central to enabling next-generation solid-state batteries.
Piezo Energy Harvesters.
Transparent Conductors.
Electrically Insulating Dielectrics.
Photovoltaic Absorbers.
Photovoltaic materials must absorb sunlight efficiently while remaining stable, earth-abundant, and non-toxic. Target absorbers have optimal band gaps (1.11.11.61.6 eV) for solar conversion, formation energies $\leq-0.5$ eV/atom, and must exclude rare or hazardous elements. This task reflects real-world sustainability constraints, forcing models to move beyond theoretical optima and identify candidates suitable for large-scale, affordable solar deployment.
Hard Coating Materials.
Hard, Stiff Ceramics.
Structural Materials for Aerospace.
Acousto-Optic Hybrids.
Materials for acousto-optic devices must balance piezoelectric and dielectric properties to minimize loading while enabling efficient coupling. Candidates are required to exhibit piezoelectric coefficients in the range $2\leq d_{ij}\leq 9$ pC/N and dielectric constants in the range $8\leq\kappa\leq 95$, with a preference for a narrow $\kappa$ band to reduce dielectric loading. Ranking emphasizes proximity to target $d$-bands and mid-range $\kappa$ values, highlighting the trade-off between acoustic response and dielectric stability.
Low-Density Structural Materials.
Toxic-Free Perovskite Oxides.
Appendix C Baselines
CDVAE.
CDVAE (Xie et al., 2021) is a conditional variational autoencoder tailored for crystal structure generation. It learns latent representations of crystals conditioned on composition, enabling the generation of valid and diverse candidate structures. We implement CDVAE using the official open-source repository [^1] https://github.com/txie-93/cdvae with default parameters, outputting the generated structures in CIF (Crystallographic Information File) format, along with latent embeddings stored as NumPy arrays for downstream analysis.
G-SchNet.
G-SchNet (Gebauer et al., 2019) is a graph-based deep generative model for molecular and crystal structures, built on the SchNet architecture. It incrementally generates atom types and positions conditioned on the partially built structure, allowing it to capture geometric and chemical validity. We implement G-SchNet using the open-source codebase [^2] https://github.com/atomistic-machine-learning/G-SchNet with default hyperparameters for $10000$ iterations. The generated samples are stored in a database, which is subsequently converted into candidate-specific CIF files.
DiffCSP.
End2end.
We implemented a GPT-4o-mini to generate candidates conditioned on the design task and its corresponding property constraints. The LLM operated with a sampling temperature of $\tau=0.8$ to promote diversity while preserving structural coherence. It was explicitly instructed to output candidates in Crystallographic Information File (CIF) format, ensuring standardized structural representations suitable for subsequent validation and property evaluation using the surrogate-assisted oracle prediction.
LLMatDesign.
LLMatDesign (Jia et al., 2024) leverages LLMs for material design by prompting LLMs iteratively improve the provided material to return the material with the single target property. As LLMatDesign is primarily designed for single-objective optimization, we adapt its released implementation [^4] https://github.com/Fung-Lab/LLMatDesign for our multi-objective materials discovery benchmarks. For candidate generation, we adopt their original prompt template, which provides a material’s chemical formula and task-specific properties and asks the model to propose a modification that satisfies the constraints. The model selects one of four modification types—exchange, substitute, remove, or add—and outputs both the modification and a natural-language hypothesis justifying the change.
Appendix D LLEMA
D.1 Implementation Details
Material Design.
Figure 7 illustrates an example prompt for the Wide-Bandgap Semiconductor task. The prompt begins with general instructions specifying the LLM’s role and objective, followed by task-specific details such as the description of the design problem, property constraints (e.g., required band gap and formation energy), and a set of previously explored candidates with their associated property values sampled from the experience buffer. We then sample $b=2$ candidate outputs at a temperature of $\tau=0.8$, chosen to balance creativity with adherence to constraints while exploiting promising directions in the search space. To further ensure physical plausibility, we introduce rule-based constraints during the evolutionary phase (after the initial generation step, $n=0$), which guide the LLM through the material discovery cycle and promote chemically valid candidates. By providing task descriptions, prior evaluations, and rule-based constraints in the prompt, we effectively steer the LLM toward feasible, property-aligned material candidates.
Crystallographic Representation.
Data-Driven Evaluation.
Figure 7: Example of input prompt for the wide band semiconductors task, including task specification, evolution constraints, in-context demonstrations, and CIF generation.
Experience Management.
We adopt an island-based evolutionary strategy (Romera-Paredes et al., 2024; Shojaee et al., 2024; Abhyankar et al., 2025) to manage the experience buffer, where generated material candidates and their evaluation scores are distributed across $m=5$ independently evolving islands. Each island is initialized with a small set of seed candidates sampled from the Materials Project API and refined using the LLM. Within each island, the buffer is divided into two components: a reward buffer, which stores candidates that fully satisfy all task-specific constraints (e.g., band gap and formation energy), and an error buffer, which stores candidates that only partially satisfy or completely violate the constraints. This separation enables the framework to reinforce high-quality generations while also retaining failure cases, which serve as negative examples to guide exploration. The experience buffer is further leveraged to construct prompts for subsequent LLM calls. After the prompt template is updated with task-specific information, one of the $m$ islands is selected at random, and $k=2$ candidates are sampled from its buffers to serve as in-context demonstrations. Candidate selection follows a Boltzmann strategy (De La Maza & Tidor, 1992) that assigns higher probability to clusters with stronger evaluation scores. Specifically, if $s_{i}$ denotes the score of the $i$-th cluster, the probability $P_{i}$ of selecting it is given by:
$P_{i}=\frac{\exp(s_{i}/\tau_{c})}{\sum_{i}\exp(s_{i}/\tau_{c})},\qquad\tau_{c}=T_{0}\left(1-\frac{u\bmod N}{N}\right),$
where $\tau_{c}$ is the temperature parameter, $u$ is the current number of candidates on the island, and $T_{0}=0.1$ and N=10, [^0] N=10{,}000 are hyperparameters. Once a cluster is selected, we sample candidates from it for inclusion in the next generation.
This mechanism integrates information from both successful and failed generations while preserving diversity across islands, thereby guiding the LLM toward more effective material discovery.
D.2 Additional Details
Surrogate Model.
@@ D.2 Additional Details
The oracle in LLEMA is designed to provide scalable and reliable property estimation by combining external databases with pretrained surrogates. After generating crystallographic representations (CIFs) for each candidate, we first query the Materials Project API [^5] https://next-gen.materialsproject.org/api to retrieve available properties. However, the API has limited coverage: many target properties such as electrical conductivity and dielectric constants are either missing for several materials or absent altogether, and the database itself, though large, cannot cover all generated candidates. To overcome this limitation, we conducted preliminary experiments with several pretrained models and identified ALIGNN (Choudhary & DeCost, 2021) and CGCNN (Xie & Grossman, 2018) as the most reliable surrogates across key properties. We use ALIGNN checkpoints trained on the JARVIS-DFT dataset [^6] https://figshare.com/articles/dataset/ALIGNN_models_on_JARVIS-DFT_dataset/17005681/6 via their official implementation [^7] https://github.com/usnistgov/alignn and the official CGCNN release [^8] https://github.com/txie-93/cgcnn. This design ensures a clear mapping: when properties are available in the Materials Project, we use them directly; when they are not, the most appropriate surrogate model is selected based on our prior experimentation. By integrating database queries with pretrained ML predictors, the oracle balances accuracy, scalability, and coverage, enabling consistent evaluation across all discovery tasks. Table 4 provides a mapping between the surrogate models and the source pretrained files used to predict the corresponding physicochemical properties.
@@
LLEMA uses a multi-objective scoring function to assign scores to the proposed candidates to guide the evolution process. The multi-objective scoring function in LLEMA is defined as:
$S(\mathcal{T,C};\mathcal{M}_{j})=\sum_{i=1}^{k}w_{i}\cdot\Phi_{i}(f_{i}(\mathcal{M}_{j}),c_{i}),$
@@
where $f_{i}(\mathcal{M}_{j})$ denotes the predicted value of the $i$-th property for candidate $\mathcal{M}_{j}$, and $c_{i}$ is the corresponding design constraint. The function $\Phi_{i}(\cdot,\cdot)$ quantifies the degree of satisfaction between the predicted property and its target constraint, normalized to a bounded reward space. The weighting coefficients $w_{i}$ encode the task-specific importance of each property, allowing LLEMA to balance competing objectives during optimization. However, for simplicity, we assign equal weights to all the properties associated with a task. For example, in the wide-bandgap semiconductor task, band gap and formation energy are given equal priorities, while energy-above-hull are included with lower but non-zero weights to ensure stability and synthesizability. Material constraints are typically of three types with better materials being higher the value being better, or lower the value being better. for example in a wide-bandgap semiconductors, band gap is supposed to be higher than $2.5$ eV whereas for formation energy, it is better to have lower values i.e. less than 1-1 eV/atom. Similarly, there are value or range based constraints Across all tasks, these weights are curated in consultation with domain heuristics: performance-critical properties are emphasized, while feasibility constraints (formation energy, density, hull stability) act as secondary filters to prevent chemically implausible candidates. This principled weighting ensures that multi-objective optimization reflects the physical and industrial priorities of each benchmark task, rather than being tuned arbitrarily.
Evolutionary Generation Rules.

### v2

@@ 4.3 Impact of Surrogate Models
Figure 5: Hit-Rate and Stability performance of LLEMA with and without surrogate model-based property prediction.
Figure 5 quantifies the effect of surrogate model–based property prediction on LLEMA’s performance. All experiments were run for 250 iterations with task constraints relaxed to encourage exploration, similar to Section 4.1. To isolate the contribution of ML-based surrogates, we removed surrogate ML models like CGCNN and ALIGNN, and restricted the workflow to using only the Materials Project database for property annotations. Furthermore, we experimented with fewer iterations while relaxing task constraints for the wide-bandgap semiconductors dataset. Even in this setting, both hit-rate and stability collapse to near-zero (<5<5%), as the evolutionary process cannot assign meaningful rewards without surrogate models to candidates for missing property annotations in the Materials Project API. The search then drifts toward trivial or repeated compounds instead of progressing toward novel solutions. By contrast, reintroducing surrogate predictors produces more than a sixfold improvement, increasing hit-rate and stability into the 25–30% range. The surrogate estimates for the out-of-distribution candidates provide the evaluation signals needed to sustain exploration beyond the sparse coverage of existing datasets. These results demonstrate that surrogate models are indispensable for providing informative fitness signals under sparse supervision, preventing collapse, and enabling effective discovery.
@@ 5 Related Work
Materials discovery has progressed from trial-and-error (Nagamatsu et al., 2001) to ab initio modeling (Jain et al., 2011; Hautier et al., 2012; Pyzer-Knapp et al., 2015), with DFT and high-throughput screening as standard tools. Machine learning further accelerates discovery via rapid property prediction (Xie and Grossman, 2018; Chen et al., 2019; Choudhary and DeCost, 2021) and generative design (Gebauer et al., 2019; Xie et al., 2022; Jiao et al., 2023), including multi-objective optimization (Gopakumar et al., 2018; Jablonka et al., 2021). Yet, these methods remain limited by data scarcity and poor transferability across domains. In contrast, LLMs have broad scientific knowledge that enables reasoning in data-poor regimes, making them a natural foundation for knowledge-guided materials design frameworks.
@@
Recent advances in generative models have shown their ability to generalize across diverse tasks using pre-trained knowledge and simple prompting strategies (Brown et al., 2020; Reddy and Shojaee, 2025; Wei et al., 2022). However, their outputs are often unreliable or inconsistent (Madaan et al., 2024; Zhu et al., 2023), motivating the use of evolutionary optimization frameworks where LLMs act as generators and external evaluators guide selection and refinement (Lange et al., 2024; Lehman et al., 2023; Liu et al., 2024; Zheng et al., 2023). Such frameworks have been successfully applied in areas including code and prompt generation (Guo et al., 2024), scientific discovery (Shojaee et al., 2025a; b), mathematical optimization (Yang et al., 2024b), program synthesis (Romera-Paredes et al., 2024), robotics reward design (Ma et al., 2024), feature engineering (Abhyankar et al., 2025), and chemical discovery (Wang et al., 2025). We extend this line of work to real-world materials science challenges by incorporating chemistry-informed LLM evolution that enforces structural validity and physical plausibility, enabling principled discovery under multi-objective constraints.
@@
Early work applied LLM-based frameworks to literature mining, named-entity recognition, and property extraction (Gupta et al., 2022; 2024; Niyongabo Rubungo et al., 2025). Beyond text extraction, more recent methods use LLMs for hypothesis generation (Miret and Krishnan, 2024), synthesis route planning, and as reasoning engines in multi-agent systems (Zhang et al., 2024; Kang and Kim, 2024; Kumbhar et al., 2025). Systems such as MatAgent (Bazgir et al., 2025) and LLMatDesign (Jia et al., 2024) combine LLM reasoning with property predictors and optimization loops, but often emphasize narrow objectives and weak constraint enforcement, leading to an unguided or infeasible search. Closely related efforts, such as MatLLMSearch (Gan et al., 2025), use LLM proposals in an evolutionary refinement loop for crystal discovery, and MatterGen (Zeni et al., 2025) uses a diffusion-based model with conditional sampling for the inverse design. Although effective, these approaches rely on either learned generative priors or limited objective formulations. LLEMA instead frames the design of materials as a multi-objective search problem with synthesizability, integrating LLM reasoning with surrogate predictors, domain constraints, and memory-based refinement to balance novelty, feasibility and property alignment.
@@ Acknowledgements
This research was partially supported by the U.S. National Science Foundation (NSF) under Grant No. 2416728 and Autodesk Research. The authors thank Modal for providing computational resources that supported the hosting and implementation of the models used in this study. Saaketh Desai is supported in part by the Center for Integrated Nanotechnologies, an Office of Science user facility operated for the U.S. Department of Energy. This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all rights, title, and interest in and to the article and is solely responsible for its contents. The United States Government retains, and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan.
@@ Appendix A Autonomous Materials Discovery
The discovery of novel materials with tailored properties is fundamental to technological progress, contributing to a global materials industry generating approximately \$ 8 trillion in revenue in 2023 (McKinsey, 2024). The impact of tailored materials is particularly significant in critical domains such as energy storage Ling (2022), photovoltaics Solak and Irmak (2023), and microelectronics Refai-Ahmed et al. (2024), where advanced materials can enhance efficiency, reduce energy footprint, and improve sustainability. The grand challenge in materials discovery lies in navigating an enormous design space encompassing a vast number of material compositions and manufacturing techniques while achieving the target material structures and properties across multiple length and time scales (Oganov et al., 2019). Human intuition-based methods, along with traditional design of experiments and computational investigations, are slow and ineffective at screening this vast space of possible materials and synthesis conditions. Consequently, the field has witnessed a rapid emergence of data-driven discovery paradigms Agrawal and Choudhary (2016) and autonomous self-driving laboratories Abolhasani and Kumacheva (2023), integrating artificial intelligence, robotics, and high-throughput computation. Data-centric platforms such as the Materials Project (Jain et al., 2013) exemplify this shift, offering open high-throughput DFT databases that enable large-scale screening of candidate materials (Horton et al., 2025). Agent-based frameworks further advance this paradigm by autonomously navigating complex chemical spaces and identifying novel stable compounds (Montoya et al., 2020). Together, these systems mark a paradigm shift, transforming materials discovery from a process of human-guided trial and error to one of algorithmic intuition and autonomous exploration.
Appendix B Additional Details
B.1 Surrogate Model
@@
The oracle in LLEMA is designed to provide scalable and reliable property estimation by combining external databases with pretrained surrogates. After generating crystallographic representations (CIFs) for each candidate, we first query the Materials Project API [^1] https://next-gen.materialsproject.org/api to retrieve available properties. However, the API has limited coverage: many target properties, such as electrical conductivity and dielectric constants, are either missing for several materials or absent altogether, and the database itself, though large, cannot cover all generated candidates. To overcome this limitation, we conducted preliminary experiments with several pretrained models and identified ALIGNN (Choudhary and DeCost, 2021) and CGCNN (Xie and Grossman, 2018) as the most reliable surrogates across key properties. We use ALIGNN checkpoints trained on the JARVIS-DFT dataset [^2] https://figshare.com/articles/dataset/ALIGNN_models_on_JARVIS-DFT_dataset/17005681/6 via their official implementation [^3] https://github.com/usnistgov/alignn and the official CGCNN release [^4] https://github.com/txie-93/cgcnn. This design ensures a clear mapping: when properties are available in the Materials Project, we use them directly; when they are not, the most appropriate surrogate model is selected based on our prior experimentation. By integrating database queries with pretrained ML predictors, the oracle balances accuracy, scalability, and coverage, enabling consistent evaluation across all discovery tasks. Table 5 provides a mapping between the surrogate models and the source pretrained files used to predict the corresponding physicochemical properties.
@@ B.2 Fitness Assessment
LLEMA uses a multi-objective scoring function to evaluate candidate materials and guide the evolutionary search process. The scoring function is defined as:
$S(\mathcal{T},\mathcal{C};\mathcal{M}_{j})=\sum_{i=1}^{k}w_{i}\cdot\Phi_{i}\!\left(f_{i}(\mathcal{M}_{j}),c_{i}\right),$
where $f_{i}(\mathcal{M}_{j})$ denotes the predicted value of the $i$-th property for candidate $\mathcal{M}_{j}$, $c_{i}$ is the corresponding design constraint, and $w_{i}\geq 0$ is a weighting coefficient that encodes the relative importance of property $i$. The function $\Phi_{i}(\cdot,\cdot)$ measures the degree of satisfaction between a predicted property value and its associated constraint, and maps it to a bounded reward space. A candidate is considered feasible if and only if $\Phi_{i}\geq 0$ for all constraints $i$. Each constraint $c_{i}$ is assumed to be one of three types: (a) a lower-bound constraint $\big(f_{i}(\mathcal{M}_{j})\geq l_{i}\big)$, (b) an upper-bound constraint $\big(f_{i}(\mathcal{M}_{j})\leq u_{i}\big)$, or (c) an interval constraint $\big(f_{i}(\mathcal{M}_{j})\in[l_{i},u_{i}]\big)$. The normalized reward $\Phi_{i}$ is defined as a signed, clipped feasibility margin that depends only on the constraint limits. Specifically, for a lower-bound constraint,
$\Phi_{i}=\mathrm{clip}\!\left(\frac{f_{i}(\mathcal{M}_{j})-l_{i}}{\max(|l_{i}|,1)},-1,1\right),$
for an upper-bound constraint,
$\Phi_{i}=\mathrm{clip}\!\left(\frac{u_{i}-f_{i}(\mathcal{M}_{j})}{\max(|u_{i}|,1)},-1,1\right),$
and for an interval constraint,
$\Phi_{i}=\mathrm{clip}\!\left(\frac{\min\!\big(f_{i}(\mathcal{M}_{j})-l_{i},\;u_{i}-f_{i}(\mathcal{M}_{j})\big)}{\max(|u_{i}-l_{i}|,1)},-1,1\right),$
@@
where $\mathrm{clip}(z,-1,1)=\max(-1,\min(1,z))$. For example, in wide-bandgap semiconductors, band gap is supposed to be higher than $2.5$ eV whereas for formation energy, it is better to have lower values i.e. less than 1-1 eV/atom and scores are calculated accordingly. The weighting coefficients $w_{i}$ encode the task-specific importance of each property, allowing LLEMA to balance competing objectives during optimization. These weights act as hyperparameters, curated in consultation with domain heuristics: performance-critical properties are emphasized, while feasibility constraints (stability-related) act as secondary filters after scoring the candidates to prevent chemically implausible candidates. However, for simplicity, we assign equal weights to all the properties associated with a task. For example, in the wide-bandgap semiconductor task, band gap and formation energy are given equal priorities, while energy-above-hull is used to ensure stability and synthesizability. This principled scoring, along with a stability check, ensures that the discovery process reflects the physical and industrial priorities of each benchmark task, rather than being tuned arbitrarily.
B.3 Multi-Island Exploration
LLEMA utilizes the multi-island framework to evolve its candidate population. At initialization, the global population is partitioned into $k$ independent islands by evenly splitting the candidate set into $k$ subpopulations. Each island maintains its own success and failure buffers and evolves independently of the others. Given a total evolutionary budget of $T$ iterations, each island receives approximately $T/k$ iterations. Therefore, the choice of $k$ determines the balance between exploration and exploitation. A smaller $k$ allows deeper refinement per island (greater exploitation), and a larger $k$ enables more independent trajectories with shallower refinement (greater exploration). Because all islands run in parallel with distinct memory states, $k$ primarily determines how the total computational budget is distributed across parallel search paths. We evaluate three representative settings: $k=1$ (single trajectory), $k=5$ (moderate multi-trajectory), and $k=10$ (highly parallel) for 250 iterations with relaxed constraints. Results for four representative tasks are shown in Table 6. Moderate island counts (for example, $k=5$) consistently provide the best balance: they allow broad exploration while still giving each island enough iterations to refine promising regions. A very small $k$ restricts exploration, and a very large $k$ reduces the per-island refinement depth. As each island maintains its own success and failure memories and adapts independently, the trajectories diverge early and explore different portions of the design space. As long as $k$ is chosen in proportion to the available compute budget (for example, using more islands only when $T$ is sufficiently large), overall performance remains stable across tasks.
Table 6: Effect of island count on representative tasks.
# Islands Wide-Bandgap High-k SAW/BAW Hard, Stiff
Semicond. Dielectrics Acoustics Ceramics
1 18.62 5.59 22.00 5.65
5 33.62 19.96 59.88 30.99
10 24.46 20.34 26.78 16.19
B.4 Evolutionary Generation Rules